From a class of 6 male and 8 female students, a committee of 4 students is chosen. How many different committees contain more than 2 female students?

I thought of doing C (6,1) * C(8,3), but not sure if it makes sense.

Answer by Edwin McCravy(20054)

From a class of 6 male and 8 female students, a committee of 4 students is chosen. How many different committees contain more than 2 female students?
That's part of the answer, the number when the committee consiste of 1 male 
and 3 females, but there is a second case!  That is when the committee consists 
of all females and no males.

Case 1.  Committees with exactly 3 females and 1 male.
         (8 females choose 3)*(6 males choose 1) = (8C3)(6C1) = (56)(6) = 336.
Case 2.  Committees with exactly 4 females and no males.
         (8 females choose 4)= (8C4)= 70.

Grand total: 336+70 = 406 ways.

Edwin